1. Fill in the blank to make the smallest number:
____ ____ × ____ = ____
2. Fill in the blank to make the biggest number:
____ ____ × ____ = ____

Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

1. Fill in the blank to make the smallest number:
____ ____ × ____ = ____
-9*9 = -81
====================
2. Fill in the blank to make the biggest number:
____ ____ × ____ = ____
99*9 = 891

Answer by Edwin McCravy(20055)

I suppose you mean a 2-digit number times a 1-digit number to
give a 1 digit number.  Then you may pick any 2-digit number 
for the first two blanks but you must then multiply it by the 1-digit 
number 0 to get 0.  0 is both the smallest and largest, in fact, the
ONLY 1-digit number you can multiply a two-digit number by to get
a one-digit number, and that 1-digit number can only be 0.

10×0 = 0
99×0 = 0

Edwin